if cause a minus b is equal to one by two and sin a + b is equal to 1 by 2 find the smallest positive value of a and b and also their most General valuesHow to find its general term??

Grade:11

1 Answers

RAMCHANDRARAO
7 years ago
a-b=0.5    sin(a+b)=0.5 or sin(a+b)=sin(30) that implies a+b=30
solve the two equations 2a=61/2 therefore a=61/4 then substitute a value into equation1
therefore b=59/4